# Mantis: Interleaved Multi-Image Instruction Tuning

### v1

@@ 3.2 Evaluation Benchmarks
For multi-image vision-language tasks, we use 2 held-in benchmarks: NLVR2 [24] and Qbench [30]; and 3 held-out benchmarks: Mantis-Eval, BLINK [18], and MVBench [19]. We report detailed statistics of different datasets in  Table 2.
NLVR2 [24] evaluates the model’s abi.lity to conduct logical reasoning across the contents of images. It asks the model to compare the contents of the two given images and judge whether a given statement is correct or not. All questions are multiple-choice. We use the test-public split for evaluation
@@
LLaVA-1.5 7B merge 53.88 49.32 31.34 37.13 36.00 41.5
LLaVA-1.6 7B merge 58.88 54.80 45.62 39.55 40.90 47.9
@@
OpenFlamingo 9B sequence 36.41 19.60 12.44 39.18 7.90 23.1
Idefics1 9B sequence 54.63 30.60 28.11 24.69 26.42 32.9
Emu2-Chat 37B sequence 58.16 50.05 37.79 36.20 39.72 44.4
Mantis-Fuyu 8B sequence 72.41 54.10 37.79 39.87 39.12 48.6
Mantis-CLIP 8B sequence 84.66 66.00 55.76 47.06 48.30 60.3
Mantis-SigLIP 8B sequence 87.43 69.90 59.45 46.35 50.15 62.6
$\Delta$ over SoTA - - +0.56 +12.90 +9.68 +1.88 +8.00 +9.1
@@ 3.3 Results on Multi-Image Tasks
InstructBLIP 7B 33.6 75.2 38.3 30.6 45.2 70.6 24.4 48.9
[…]
LLaVA-1.5 7B 58.2 76.6 64.8 35.3 53.4 70.4 25.6 59.8
Qwen-VL-Chat 7B 63.8 78.2 61.8 35.9 56.6 68.2 15.5 60.8
[…]
Idefics1 9B 39.3 68.8 45.3 32.5 50.4 51.6 21.1 48.0
OpenFlamingo 9B 46.3 58.0 32.4 28.7 51.4 45.7 18.6 43.8
@@
We report our results on the multi-image task in  Table 3. We put the GPT-4V’s results at the top of the table as it’s a closed-source LMM. We put comparison results between Mantis with other powerful open-source LMMs, along with the insights derived in the following paragraphs.
Mantis performs well on held-in evaluation Mantis-8B-SigLIP achieves $87.43$ and $69.90$ on NLVR2 and Q-Bench respectively, marking it the best open-source LMM on these 2 tasks. When compared to Idefics2, which also trains on NLVR2 training set, our result is marginally better. It’s worth noticing that the Mantis-8B-SigLIP is only 1.37 behind GPT-4V on NLVR2. The highest scores on these 2 benchmarks demonstrate that Mantis has gained strong ability in multi-image logical reasoning and low-level vision understanding.
Mantis generalizes well on held-out evaluation Generalization of held-out multi-image tasks is a more valuable metric. Our evaluation results on the 3 held-out benchmarks, Mantis-Eval, BLINK (val), and MVBench also demonstrate our strong generalization ability. Mantis-8B-SigLIP achieves $59.45$, $46.35$, and $50.15$ respectively on the 3 held-out tasks, surpassing all other open-source baselines. The gain over existing strongest baseline are particulouly tremendous on Mantis-Eval and MVBench.
Decoder-only LMM vs encoder-decoder LMM Unlike many other LMMs like Flamingo, LLaVA, and BLIP, Fuyu is a decoder-only LMM without a vision encoder. A big advantage of this practice is that it supports super-resolution, so the image will not be resized to fit the vision encoder. So which architecture is better? To investigate deeper, we also train an additional Mantis-8B-Fuyu as a comparison. Results in  Table 3 show that the performance of Mantis-8B-Fuyu is worse than Mantis-8B-CLIP and Mantis-8B-SigLIP where a vision encoder exists, even after fine-tuning on the same set of data. We thus conclude vision encoder should be necessary. This is also consistent with our intuition, where people see more, but think/write/talk less. In short, Encoder-Decoder LMM might be a better choice for LMM architecture.
Vision encoder matters As stated above, the Encoder-Decoder is a superior architecture compared to the decoder-only LMM. Following this, the quality of the vision encoder can also have big effects on the final LMM performance. We include 2 versions of Mantis, which is trained with CLIP [22] and SigLIP [23] respectively. And Mantis-8B-SIGIP can achieve generally better performance than the Mantis-8B-CLIP. This shows that SigLIP is a better vision encoder for LMM development.
‘Merge’ vs ‘Sequence’ input The ‘merge’ evaluation is an important baseline to evaluate the multi-image understanding ability of those LMMs that can accept one image at a time. If the results are good by merging images, then interleaved text/image instruction tuning will seem unnecessary. However, it’s not the truth. LLaVA-v1.6’s dynamic image resolution feature can divide a single image into multiple tiles, processing like multiple images [7]. But it still behaves badly on multi-image tasks, only getting $45.62$ and $39.55$ on the Mantis-Eval and BLINK benchmark. Though the results are better than LLaVA-v1.5, which does not have the dynamic image resolution feature, it’s still limited by the image resolution and cannot recognize separate images well.
@@ 3.4 Results on Single-Image Tasks
We also evaluate Mantis-8B-CLIP and Mantis-8B-SigLIP on various single-image tasks, including TextVQA [48], VQA-v2 [49], MMBench [50], MMMU [51], etc. Results are shown in  Table 4. All the evaluations are conducted with the help of LMMs-Eval [52] tool.
Compared to a set of previous famous LMMs, Mantis-8B-SigLIP gets significant improvements on MMBench-English, MMMU, and ScienceQA benchmarks. Solving problems in these tasks requires not only the visual perception ability but also the good reasoning ability of the LLM backbone. Considering that we are using the powerful LLaMA-3 as the backbone LLM, these improvements are thus expected. However, it’s also worth noticing the performance degeneration on some tasks after introducing the multi-image ability. Compared to LLaVA-v1.5-7B, Mantis-8B-CLIP’s performance on TextVQA, VQA-v2 drops $1.8$ and $3.6$ respectively. However, this decrease is acceptable as a trade-off for enhancing multi-image ability.
@@ 3.5 Case Studies
Figure 3: Case Study of Mantis. The Mantis model is able to understand multiple images and complete the given task
As shown in  Figure 3, we present 2 cases where Mantis is doing do compared to LLaVA-1.5, which does not have multi-image ability, along with Emu-2 and GPT-4V, which also have multi-image ability. Mantis can understand multiple images and generate a more accurate answer.
@@ 4.1 Large Multimodal Models
Large Multimodal Models (LMMs) are mostly denoted as large language models that can understand multiple modalities beside the human language. Some works aim to fuse image, audio, video, or more modalities with language [53, 54], while more works mainly focus on better combining vision knowledge with the language [55, 14, 45]. BLIP-2 has curated a large-scale image captioning dataset, bootstrapping a language model along with a vision encoder to be a powerful multimodal model [8]. Following this, LLaVA further proposes a cheaper approach to train a powerful LMM through visual instruction tuning [32]. LLaVA-Next further optimizes single-image performance, with a cost of an increasing number of image tokens per token [7], consuming more than 2k for each image, which is about 4 times of the original LLaVA model. Later LMMs like QwenVL [31], CogVLM [9], Yi-VL [47], etc. all follow a similar architecture of LLaVA. However, real-world visual problems are way more complex and usually require a sequence of images to describe the situation. Although better performance is achieved in the single-image scenario, it’s doubtful whether it’s worth the cost.
@@ 4.3 Multi-Image Evaluation
While many benchmarks have been developed for single-image tasks, such as TextVQA [48], VQA-V2 [49], MMBench [50], and so on, the multi-image evaluation did not capture much of the people’s attention until recently. Difference description is the first task which large-scale data and good evaluation benchmarks are developed for, including works like Spot-the-Diff [25], Birds-to-Words [27], etc. They are usually evaluated by n-gram metrics like BLEU, and ROUGE, and only involve two images. The later NLVR2 [24] further includes logical reasoning across 2 images into evaluation, focusing on a higher-level of cognition ability. Q-bench [30] and BLINK [18] are both proposed for low-level multi-image vision tasks, where human easily do but model struggles. Another set of multi-image evaluation comes from video or image sequence evaluation, including Mementos [17], MVBench [19], etc. In this case, model usually need to take more than 2 images for answering.
@@ Appendix A Appendix / supplemental material
A.1 Heuristics for data curation
@@
Appendix B Prompt Template

### v2

@@ 3.2 Evaluation Benchmarks
For multi-image tasks, we use 2 held-in benchmarks: NLVR2 [18] and Qbench [28]; and 3 held-out benchmarks: Mantis-Eval, BLINK [29], and MVBench [30]. We report detailed statistics in  Table 3.
NLVR2 [18] evaluates the model’s ability to conduct logical reasoning across the contents of images. It asks the model to compare the contents of the two given images and judge whether a given statement is correct or not. All questions are multiple-choice. We use the test-public split for evaluation.
@@
LLaVA-v1.5 7B merge 53.88 49.32 31.34 37.13 36.00 41.5
LLava-V1.6 7B merge 58.88 54.80 45.62 39.55 40.90 48.0
@@
BLIP-2 13B merge 59.42 51.20 49.77 39.45 31.40 46.2
InstructBLIP 13B merge 60.26 44.30 45.62 42.24 32.50 45.0
CogVLM 17B merge 58.58 53.20 45.16 41.54 37.30 47.2
OpenFlamingo 9B sequence 36.41 19.60 12.44 39.18 7.90 23.1
Otter-Image 9B sequence 49.15 17.50 14.29 36.26 15.30 26.5
Idefics1 9B sequence 54.63 30.60 28.11 24.69 26.42 32.9
VideoLLaVA 7B sequence 56.48 45.70 35.94 38.92 44.30 44.3
Emu2-Chat 37B sequence 58.16 50.05 37.79 36.20 39.72 44.4
@@
Mantis-CLIP 8B sequence 84.66 66.00 55.76 47.06 48.30 60.4
Mantis-SigLIP 8B sequence 87.43 69.90 59.45 46.35 50.15 62.7
Mantis-Flamingo 9B sequence 52.96 46.80 32.72 38.00 40.83 42.3
Mantis-Idefics2 8B sequence 89.71 75.20 57.14 49.05 51.38 64.5
$\Delta$ over SOTA - - +2.84 +18.20 +8.30 +3.87 +1.98 +11.0
@@ 3.3 Results on Multi-Image Tasks
OpenFlamingo 9B 46.3 58.0 32.4 28.7 51.4 45.7 18.6 40.2
Idefics1 9B 39.3 68.8 45.3 32.5 50.4 51.6 21.1 44.1
InstructBLIP 7B 33.6 75.2 38.3 30.6 45.2 70.6 24.4 45.4
Yi-VL 6B 44.8 72.5 68.4 39.1 51.3 71.7 29.7 53.9
Qwen-VL-Chat 7B 63.8 78.2 61.8 35.9 56.6 68.2 15.5 54.3
LLaVA-1.5 7B 58.2 76.6 64.8 35.3 53.4 70.4 25.6 54.9
@@
Mantis-CLIP 8B 56.4 73.0 66.0 38.1 53.0 73.8 31.7 56.0
Mantis-SigLIP 8B 59.2 74.9 68.7 40.1 55.4 74.9 34.4 58.2
[…]
We report our results on the multi-image task in  Table 4. We put the GPT-4V’s results at the top of the table as it’s a closed-source LMM. We can see that our best version Mantis-Idefics2 already matches the overall performance of GPT-4V. We put comparison results between Mantis and other open-source LMMs, along with the insights derived in the following paragraphs.
Mantis performs well on held-in evaluation We found that Mantis-SigLIP without any multi-image pre-training can already achieve best known performance on these two benchmarks. Mantis-Idefics2 has better performance of $89.71$ and $75.20$ on NLVR2 and Q-Bench respectively, marking it the SoTA for these two tasks. Notably, Mantis-Idefics2 also beats the Idefics2 by $2.84$ points, which was already trained on NLVR2. It shows the cross-task benefits of our instruction tuning. The remarkable performance on these 2 benchmarks demonstrate that Mantis has gained strong ability in multi-image logical reasoning and low-level vision understanding.
Mantis generalizes well on held-out evaluation We show that our models can attain strong performance on other held-out benchmarks, Mantis-Eval, BLINK (val), and MVBench. Mantis-SigLIP achieves $59.45$ on Mantis-Eval, while Mantis-Idefics2 achieves $49.05$, and $51.38$ on BLINK and MVBench respectively, surpassing all other baselines on these held-out tasks. The gain over the existing strongest baseline is particularly tremendous on Mantis-Eval and MVBench.
Multi-Image Pre-training is not necessary In these experiments, Mantis-CLIP and Mantis-SigLIP are attaining much better performance than Mantis-Flamingo and similar performance as Mantis-Idefics. These results reveal that the multi-image pre-training (on massive web corpus) is not a necessary path toward strong multi-image performance.
‘Merge’ vs ‘Sequence’ input The ‘merge’ evaluation is an important baseline to evaluate the multi-image understanding ability of those LMMs that can only accept one image at a time. LLaVA-v1.6’s dynamic image resolution feature can divide a single image into multiple tiles, processing the ‘merge’ image like multiple images [7]. But it still performs poorly on multi-image tasks, only getting $45.62$ and $39.55$ on the Mantis-Eval and BLINK benchmark, which are at least 10 absolute points lower than Mantis. This result indicates the advantages of building models to accept sequence image inputs.
@@ 3.4 Results on Single-Image Tasks
We also evaluate Mantis-CLIP and Mantis-SigLIP on various single-image tasks, including TextVQA [41], VQA-v2 [42], MMBench [43], MMMU [44], etc. Results are shown in  Table 5. All the evaluations are conducted with the help of LMMs-Eval [45] tool.
Compared to a set of popular LMMs, Mantis-SigLIP gets significant improvements on MMBench-English, MMMU, and ScienceQA benchmarks though we did not specifically optimize the single-image abilities. Solving problems in these tasks requires not only the visual perception ability but also the good reasoning ability of the LLM backbone. Considering that we are using the powerful LLaMA-3 as the backbone LLM, these improvements are expected. We also compare the results of Mantis-Idefics2 against Idefics2 on single-image tasks. On several tasks like MMB, MMMU, and OKVQA, Mantis-Idefics maintains a similarly strong performance as Idefics2. However, we do observe some drops in ScienceQA, MathVista, which are mainly due to a lower mix ratio of OCR and Math data in our instruction tuning. Overall, the average drop of 4% is in the tolerable range.
3.5 Ablation Studies
Is multi-image instruction tuning necessary? To investigate, we further train a Mantis-Flamingo based on OpenFlamingo [14], which has been trained on MMC4 [46], a text-image interleaved pre-training dataset. As shown in Table 4, the average performance of Mantis-Flamingo increased by $19.2$ points after further fine-tuning on Mantis-Instruct. The results show that there is still a huge improvement space, even if a model is fully pre-trained on a multi-image dataset, which demonstrates the necessity of learning multi-image skills in the fine-tuning phase.
Is multi-image pre-training necessary? To investigate, we design an ablation study, where one model is pre-trained on the CC3M 560K subset, which contains only image-text pairs, while the other one is pre-trained additionally on OBELICS-100K subset, which contains interleaved multi-image-text pairs. Due to limited computation resources, we fine-tune both models on a downsampled Mantis-Instruct dataset to compare their performance. Results in  Table 6 show that there are little improvements by adding additional pre-training on OBELICS-100K. We thus conclude that multi-image pre-training could be useful, but it is not a necessary path for models to gain multi-image abilities. The poor performance on Mantis-Flamingo also confirm our hypothesis.
Table 6: Ablation on whether multi-image pre-training is necessary. One model is pre-trained on the CC3M subset, while the other is additionally pre-trained on the a OBELICS-100K multi-image dataset. They are both fine-tuned on a subset Mantis-Instruct for comparison. Results show minor improvement with additional multi-image pre-training.
Pretrain NLVR2 Q-Bench Mantis-Eval BLINK MVBench Avg
CC3M subset + MantisInstruct subset 76.73 57.00 47.00 43.79 47.08 54.32
[…]
CC3M subset + MantisInstruct subset
+OBELICS-100K
76.83 60.50 53.00 45.20 45.97 56.30
$\Delta$ (with OBELICS-100K) +0.10 +3.50 +6.00 +1.41 -1.11 +1.98
Table 7: Data ablation study results of different subsets of Mantis-Instruct, where each skill corresponding to a specific subsets defined in Table 1. $\Delta$ column of each row represents the performance improvements after adding new subsets to the last column results.
Models NLVR2 Q-Bench Mantis-Eval BLINK MVBench Avg $\Delta$
Idefics2-8B 86.87 57.00 48.85 45.18 29.68 53.52 -
[…]
+Compare 85.92 69.00 53.92 48.00 48.33 61.03 +0.47
+Reason 89.62 68.50 62.21 48.31 47.93 63.31 +2.28
+Temporal 89.54 67.00 61.29 49.34 50.02 63.44 +0.12
Ablation study of Mantis-Instruct components: We conduct a data ablation study to analyze the effects of different subsets in Mantis-Instruct to enhance a model’s multi-image ability. As shown in  Table 7, As new subsets of Mantis-Instruct are added continually, the overall performance is continually improving, proving that every subset is contributing positively.
@@ 3.6 Case Studies
To conduct a qualitative analysis of Mantis models on specific multi-image scenarios, we include some case studies in  Figure 3, where we include Emu2, which can accept a sequence of images as inputs, and LLaVA-1.5, which can only accept one image at a time, and the powerful GPT-4V, which perform well on both single-image and multi-image scenario.
Figure 3: Two case studies in multi-image scenario. Mantis can understand multiple images and complete the given task, thanks to the 4 multi-image skills learned from Mantis-Instruct.
The first case requires the model to count the number of dice respectively in the two images. While previous LMMs might have learned good counting ability in the single-image scenario, most of them failed in the multi-image counting problem, including GPT-4V. The second case requires the model to involve their world knowledge to infer the mood of the character in three images. However, Emu2, and LLaVA-1.5 output similar contents, and both fall back into the simple captioning behavior. It demonstrates that these models struggle to distinguish two images as separate information. In contrast, Mantis can correctly do the task due to the existing of reasoning subset in Mantis-Instruct.
@@ 4.1 Large Multimodal Models
Large Multimodal Models (LMMs) are mostly denoted as large language models that can understand multiple modalities besides the human language. Some works aim to fuse image, audio, video, or more modalities with language [47, 48], while more works mainly focus on better-combining vision knowledge with the language [49, 14, 36]. BLIP-2 has curated a large-scale image captioning dataset, bootstrapping a language model along with a vision encoder to be a powerful multimodal model [8]. Following this, LLaVA further proposes a cheaper approach to train a powerful LMM through visual instruction tuning [32]. LLaVA-Next further optimizes single-image performance, with a cost of an increasing number of image tokens per token [7], consuming more than 2k for each image, which is about 4 times of the original LLaVA model. Later LMMs like QwenVL [31], CogVLM [9], Yi-VL [40], etc. all follow a similar architecture of LLaVA. However, real-world visual problems are way more complex and usually require a sequence of images to describe the situation. Although better performance is achieved in the single-image scenario, it’s doubtful whether it’s worth the cost.
@@ 4.3 Multi-Image Evaluation
While many benchmarks have been developed for single-image tasks, such as TextVQA [41], VQA-V2 [42], MMBench [43], and so on, the multi-image evaluation did not capture much attention until recently. Difference description is the first task which large-scale data and good evaluation benchmarks are developed for, including works like Spot-the-Diff [19], Birds-to-Words [21], etc. They are usually evaluated by n-gram metrics like BLEU, and ROUGE, and only involve two images. The later NLVR2 [18] further includes logical reasoning across 2 images into evaluation, focusing on a higher level of cognition ability. Q-bench [28] and BLINK [29] are both proposed for low-level multi-image vision tasks, where humans easily do but model struggles. Another set of multi-image evaluations comes from video or image sequence evaluation, including Mementos [51], MVBench [30], etc. In this case, the model usually needs to take more than 2 images to answer. Recently, there are also attempts to prompt LMMs to evaluate vision-language task [52], or image-generation tasks [53], which also demands good multi-image reasoning ability. The accuracies of the these tasks of different LMMs can be seen as another form of benchmark for multi-image ability.
@@ 5 Conclusion
Limitations
Our work mainly focuses on improving the multi-image ability of LMMs. Our collected dataset Mantis-Instruct reformat some existing multi-image datasets, which might still contain some noise due to the quality of the original collection. Besides, we found that most of the examples in the Mantis-Instrut tend to be multiple-choice QA questions, where the response tends to be short, resulting in the current version of Mantis models tending to output shorter responses instead of long reasoning texts and falling short in instruction-following ability in the multi-image scenario. Although we have introduced Multi-VQA subset to alleviate this issue, this problem still exists, and we will leave it to future work by introducing more long-form multi-image data. While enhancing the multi-image ability well, we find out that there is a trend of performance degeneration in the single-image performance compared to the original model Table 5. Some single-image QA datasets have been introduced to alleviate this issue, but the conflicts between multi-image and single-image ability seem to exist continually. How to balance these two abilities will also be our future direction.
Societal Impacts
Our proposed model, Mantis, and released dataset, Mantis-Instruct, can have substantial impacts on society through the application of Large Multimodal Models. LMMs with multi-image ability can analyze, and reason in real-world scenarios, helping machines, and robots, to make decisions automatically, assisting people in website browsing, travel planning based on multiple maps and pictures, etc. We believe Mantis can serve as one of the first useful open-source LMMs that make progress in these applications to make these scenarios possible and efficient.
However, there is a potential that our Mantis can generate hallucinations, make the wrong decisions, and fail to reason across complex real-world scenarios. While we have tried to improve Mantis’s accuracy and performance in instruction-following and multi-image understanding by carefully training on high-quality Mantis-Instruct, Mantis can still sometimes make mistakes, thus harming its faithfulness as a useful LMM tool. There is also the potential for misuse of our model and dataset after we make them open source. Though we have added proper licenses and terms of usage to our assets, misuse cannot be safely eliminated.
@@ Appendix A Appendix / supplemental material
A.1 Details of Mantis-Instruct subsets
We construct Mantis-Instruct from multiple publicly available datasets. Detailed statistics are shown in  Table 1. We use a similar dataset format with LLaVA’s, where each data item contains multiple images and multiple turns of QA pairs are gathered. We report the number of examples, the number of images per item, the average conversation turns, and the average length. We describe the task type of each dataset and the processing techniques used in the following:
LLaVA-665k-multi: Besides the above-mentioned multi-image reasoning tasks, we also manually reformat the LLaVA-655k dataset from LLaVA-1.5 into a "synthetic" multi-image dataset. Since each original data item consists of multiple QA pairs about a single image, we randomly merge 2 to 4 data items into a multi-image data item. Then we add proper image denotation like "For the second image, …" for each question, and shuffle all the QA pairs to form the final multi-image data item. Answering each question in this "synthetic" dataset does not require reasoning across images, but demands high image co-reference ability. Besides, it also helps avoid forgetting single-image ability in multi-image scenarios and increases instruction-following ability.
NLVR2 [18]: NLVR2 is a natural language reasoning dataset grounded in images. The task is to determine whether a statement is true or false given the visual contents of 2 images. NLVR2 covers many linguistic phenomena, such as cardinality, existential quantifiers, universal quantifiers, etc.
ImageCoDe [54]: ImageCoDe is a multimodal challenge called image retrieval from contextual descriptions, where a multimodal model is required to retrieve/select the correct image from a set of 10 minimally contrastive images. For each image, 9 similar images are retrieved from Open Images Dataset V6 [55] using CLIP encodings.
Spot-the-diff [19]: This is a dataset that requires the model to generate the difference description given 2 images. We use the processed version from MIMIC-IT [58].
Multi-VQA: This dataset contains multi-image reasoning QA pairs, where each question is generated by GPT-4 based on the captions of the images. Each question is required to involve at least two images. Images and captions are sampled from ShareGPT-4V [56]. The prompt template is in  Table 8.
Single-image-VQA: Instead of focusing only on multi-image tasks, we claim that it’s necessary to also include some single-image VQA dataset to avoid catastrophic forgetting on single-image tasks. We included some of the single image data from LLaVA-665k as well as DocVQA [61], DVQA [62], and ChartQA [63] to enhance its ability on diagrams and OCR. In practice, since the size of DVQA is too large, we only use 30k out of 200k examples for the training.
A.2 Heuristics of data curation
@@
A.3 Training Details
For Mantis-CLIP and Mantis-SigLIP, we first pre-train the multimodal projector on LLaVA pre-train data, where the learning rate is set to 1e-3, the batch size is set to 256. After pre-training, we fine-tune the 2 models on the Mantis-Instruct. For Mantis-Flamingo and Mantis-Idefics2, we directly fine-tune them on Mantis-Instruct since they have already been pre-trained on millions of data.
During the fine-tuning, we train each model on the data for 1 epoch, with a batch size of 128. The maximum context length is set to 8192. The learning rate is all set to 1e-5, except for Idefics2, where the learning rate is set to 5e-6 to better preserve its original knowledge. We set the warmup ratio to be 0.03 and use a cosine learning rate scheduler. We speed up our training and inference with Flash-attention2 [64]. We use DeepSpeed Zero-3 [65] for full fine-tuning. We apply QLoRA [66] along with DoRA [67] to more efficiently do comprehensive ablation studies under limited resources. All full fine-tuning ran on 16 A100 GPUs while the ablation study ran on 8 A100 GPUs.
A.4 Prompt Template
Table 8: Propmt template used to curate the Multi-VQA subset.